OpenStudy (anonymous):

Find the equation of the plane containing the points (3,0,0) , (0,8,0) and (0,0,6). Show each step of your process. Im stuck. please help.

ganeshie8 (ganeshie8):

u familiar wid intercepts form ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

\(\large \frac{x}{a} + \frac{y}{b} + \frac{z}{c} = 1\)

ganeshie8 (ganeshie8):

observe that, you're given x, y, z intercepts, a = 3 b = 8 c = 6 plug them in above equation